Modify X and apply Markovs inequality to upper bound P(X > 3) when X > 2 and E[X] = 2.5.
Answer #1

Let Y = X - 2, then E[Y] = E[X - 2] = E[X] - 2 = 2.5 - 2 = 0.5

As, X \ge 2 => X - 2 \ge 0 => Y \ge 0

Now, as Y is nonnegative random variable, then by Markov's inequality,

P(Y \ge a) \le E(Y) / a for any a > 0

Now,

P(X \ge 3) = P(X - 2 \ge 3 - 2) = P(Y \ge 1)

From Markov's inequality,

P(X \ge 3) = P(Y \ge 1) \le 0.5 / 1

=> P(X \ge 3) \le 0.5